# RT-GuIDE: Real-Time Gaussian splatting for Information-Driven Exploration

### v2

@@ Abstract
We propose a framework for active mapping and exploration that leverages Gaussian splatting for constructing dense maps. Further, we develop a GPU-accelerated motion planning algorithm that can exploit the Gaussian map for real-time navigation. The Gaussian map constructed onboard the robot is optimized for both photometric and geometric quality while enabling real-time situational awareness for autonomy. We show through simulation experiments that our method yields comparable Peak Signal-to-Noise Ratio (PSNR) and similar reconstruction error to state-of-the-art approaches, while being orders of magnitude faster to compute. In real-world experiments, our algorithm achieves better map quality (at least 0.8dB higher PSNR and more than 16% higher geometric reconstruction accuracy) than maps constructed by a state-of-the-art method, enabling semantic segmentation using off-the-shelf open-set models. Experiment videos and more details can be found on our project page: https://tyuezhan.github.io/RT_GuIDE/
@@ I Introduction
Active mapping is a problem of optimizing the trajectory of an autonomous robot in an unknown environment to construct an informative map in real-time. It is a critical component of numerous real-world applications such as precision agriculture [1], infrastructure inspection [2], and search and rescue [3] missions. While nearly all tasks rely on recovering accurate metric information to enable path planning, many also require more fine-grained information. Recent advances in learned map representations from the computer vision and graphics communities [4, 5] have opened up new possibilities for active mapping and exploration while maintaining both geometrically and visually accurate digital twins of the environment.
While prior work effectively solves the problem of information-driven exploration [6, 7, 8, 9, 10] or frontier-based exploration [11, 12, 13], in this work we consider the additional problem of generating radiance fields while also performing autonomous navigation. Prior work has also proposed information metrics using novel learned scene representations that are capable of high quality visual reconstruction but are incapable of running in real-time onboard a robot. To enable efficient mapping and planning in these novel representations, we consider approximation techniques for computing the information gain. Further, we consider the problem of generating high-quality maps that are capable of novel-view synthesis that can be used for downstream applications.
Fig. 1 shows the elements of our approach. We use the Gaussian splatting approach proposed in [14] to generate maps. We propose an efficient information gain metric and use a hierarchical planning framework to plan high-level navigation targets that yield maximal information in the environment and low-level paths that are dynamically feasible, collision-free and maximize the local information of the path. Our system runs real-time onboard a fully autonomous unmanned ground vehicle (UGV) to explore an unknown environment while generating a high-fidelity visual representation of the environment.
@@
    A unified framework for online mapping and planning built only on a Gaussian splatting-based map, eliminating the need for additional volumetric representations.
@@ II Related Work
To effectively construct a map of the environment, numerous map representations have been proposed in the robotics community. The most intuitive but effective volumetric representation has been widely used. Voxel-based representation could maintain information such as occupancy [15] or signed distance [16, 17]. With the recent application of semantic segmentation, semantic maps that contain actionable information have been proposed [18, 19, 20, 21, 22]. With the recent advances in learned map representations in the computer vision community, Neural Radiance Fields (NeRF) [4] and 3D Gaussian Splatting (3DGS) [5] have become popular representations for robotic motion planning. In this work, we study the problem of active mapping with the learned map representations.
The problem of exploration and active Simultaneous Localization and Mapping (SLAM) has been widely studied in the past decade. The classical exploration framework uses a model-based approach to actively navigate towards frontiers [11, 12] or waypoints that have the highest information gain [6, 7, 8, 9, 10, 23, 24]. Some recent work combines the idea of frontier exploration and the information-driven approach to further improve efficiency [25, 26, 27, 28]. However, most of the existing work developed their approaches based on classical map representations. For example, frontiers are typically defined in voxel maps, information metrics are typically associated with occupancy maps or Signed Distance Field (SDF) maps.
@@
In this work, we instead consider the active mapping problem with a learned map representation. Bayesian neural networks [29, 30] and deep ensembles [24, 31] are common approaches for estimating uncertainty in learned representations. [32, 33] use the idea of computing the uncertainty based on consecutive measurements from a moving robot in contrast to train multiple models in the traditional ensemble frameworks. Radiance field representations provide additional possibilities for estimating uncertainty through the volumetric rendering process [34]. [35] uses the difference between rendered depth and observations, and the alignment of $\alpha$-blended and median depth as measures of uncertainty.  [36] leverages Fisher information to compute pixel-wise uncertainty on rendered images. Where prior work in implicit and radiance field representations use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a map represented by Gaussians encodes physical parameters of the scene, which motivates estimating information from the Gaussian parameters directly.
@@
Prior work has also considered planning directly in radiance fields. [37] plans trajectories in a Gaussian map and uses observability coverage and reconstruction loss stored in a voxel grid as an approximation of information gain for exploration. Sim-to-real approaches leverage the rendering quality of the learnt radiance fields to train downstream tasks such as visual localization, imitation learning [38] by effectively utilizing the representation as a simulator. [39] use a liquid neural network to train imitation policies for a quadrotor robot in a Gaussian splat. [40] utilize the geometric fidelity of the representation to first map the environment and then perform trajectory optimization to plan paths in these environments. [41] use a pre-generated Gaussian map to compute safe polytopes to generate paths and also present a method to synthesize novel views with a coarse localization of the robot and refine the estimate by solving a Perspective-n-Point (PnP) problem. The probabilistic representations of free space has also been utilized for motion planning [42] where authors use uncertainties in the learned representation to provide guarantees on collision-free paths.
@@ III Problem Specification & Preliminaries
Figure 2: The proposed active mapping framework. The proposed framework contains two major components, the planning module and the mapping module. As can be seen in the figure, the Mapping module ([A]) takes in RGB, depth and pose measurements, and updates the map representation at every step and computes the utility of cuboidal regions (Sec. IV-B1). The utility of each region is then passed to the planning module which comprises the topological graph and motion primitive library ([B]). Viewpoints are sampled around regions in the map and added to the topological graph. A high-level guidance path is generated from the topological graph to viewpoints with highest utility and passed to the trajectory planner. The trajectory planner in turn attempts to plan a path to goal that maximizes information gathering (queried from the mapper). The planned trajectory is executed by the robot to get a new set of observations.
@@
$\mathop{\arg\min}\limits_{\hat{G},\ x_{1:T}}$ $\sum_{{test}\in\mathcal{T}}\mathcal{L}(y^{*}_{test},h(x^{*}_{test},\hat{G}))$ (1)
@@
$\Phi(x_{1:k},y_{1:k}):=\mathop{\arg\min}\limits_{\hat{G}}\sum_{s=1}^{k}\mathcal{L}(y_{s},h(x_{s},\hat{G})).$ (2)
Then, given the posterior estimate of the map $\hat{G}$, we find the subsequent viewpoint $x_{k+1}$ that maximizes the mutual information $\mathcal{I}(\cdot)$ between the map $\hat{G}$ and the measurement $y_{k+1}$ observed from the corresponding pose:
$\max_{x_{k+1}}\ \mathcal{I}(\hat{G};y_{k+1}|x_{k+1}).$ (3)
@@
For a given pixel $p$ in the image, its color $\mathcal{C(\cdot)}$ and depth $\mathcal{D(\cdot)}$ can be obtained as
$\mathcal{C}(p)=\sum_{i=1}^{n}c_{i}f_{i}(p)\prod_{j=1}^{i-1}(1-f_{j}(p))\$
and
$\mathcal{D}(p)=\sum_{i=1}^{n}d_{i}f_{i}(p)\prod_{j=1}^{i-1}(1-f_{j}(p)),$
where
$f_{i}(p)=\alpha_{i}\exp\left\{-\frac{||p-\mu_{2D,i}||^{2}}{2r_{2D,i}^{2}}\right\}.$
At every iteration, we render color image $\hat{y}_{c}$ and depth image $\hat{y}_{d}$. We set the function $\mathcal{L}$ (from eq. (2)) to
$\mathcal{L}$ $=\frac{1}{N_{p}}\sum_{p\in y}\left(|\hat{y}_{d}-y_{d}|+\lambda_{1}|\hat{y}_{c}-y_{c}|\right)$ (4)
$+\lambda_{2}(1-SSIM(\hat{y}_{c},y_{c}))$
@@ IV Method
Our proposed framework comprises the mapping module and the planning module, as illustrated in Fig. 2. The mapping module (Sec. IV-A) accumulates measurements and poses to generate a reconstruction of the environment and computes the uncertainty of Gaussians in the map. This is then passed to the planning module for planning guidance paths (Sec. IV-B1) and trajectories (Sec. IV-B2).
@@ IV-A Mapping & Uncertainty Estimation
We build the 3DGS mapping module upon the [14] framework. Similar to [14], we set all Gaussians to be isotropic. As in [14], we set $\lambda_{1}=0.4,\lambda_{2}=0.1$. At the beginning of each mapping iteration, we initialize new Gaussians at positions corresponding to the current color and depth measurements. The parameters of these Gaussians are refined through backpropagation. As in [5], we prune Gaussians with low opacity or large radius.
@@
Uncertainty Estimation. The objective of finding the next-best-view can then be expressed as
$x^{*}_{k+1}=\mathop{\arg\max}\limits_{x_{k+1}}\mathcal{I}(y_{k+1};\hat{G}|x_{k+1}).$ (5)
The information gain $\mathcal{I}(\cdot)$ is defined as
$\mathcal{I}$ $(y_{{k+1}};\hat{G}|x_{k+1})$ (6)
$=\mathbb{H}(y_{k+1}|x_{k+1})-\mathbb{H}(y_{k+1}|\hat{G},x_{k+1})$
@@
where $\mathbb{H}(\cdot)$ is the entropy. We make the following assumptions in this work to approximate the information gain efficiently: (a) $\hat{G}\sim\mathcal{N}(\mu_{\hat{G}},\Sigma_{\hat{G}})$; (b) $y_{k+1}|\hat{G}\sim\mathcal{N}(h(x_{k+1},\hat{G}),\Sigma_{y})$; (c) constant isotropic variance $\Sigma_{y}$. From (a) and (b), the information gain can be simplified as
$\mathcal{I}(y_{{k+1}};\hat{G}|x_{k+1})$ $=\frac{1}{2}\log\det\left(\frac{J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal}+\Sigma_{y}}{\Sigma_{y}}\right)$ (7)
$=\frac{1}{2}\log\det(I+\Sigma_{y}^{-1}J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal})$
@@
where $J_{k+1}=\frac{\partial h(x_{k+1},\hat{G})}{\partial G}$. Using the first order approximation of $\log\det$ about the identity, we can define a proxy metric for scoring next best views as
$x^{*}_{k+1}=\mathop{\arg\max}\limits_{x_{k+1}}\mathrm{Tr}(\Sigma_{y}^{-1}J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal}).$ (8)
@@
As in [36], assumption (c) allows us to simplify the optimization by ignoring the effect of $\Sigma^{-1}_{y}$. Evaluating $J_{k+1}$ is computationally expensive because it typically requires rendering at new poses. To achieve real-time evaluation of multiple candidate viewpoints, we replace the Jacobian term $J_{k+1}$ with a binary matrix that corresponds to its sparsity pattern. Intuitively, this captures the set of Gaussians within the field of view at pose $x_{k+1}$. The key insight behind this approximation is that we estimate information gain using the current uncertainty, based on the assumption that local measurements lead to precise local map estimates. We show empirically that the exploration performance based on our metric (in Sec. V-A) is similar to the performance based on eq. (8). We propose a heuristic for estimating uncertainty $\Sigma_{\hat{G}}$ based on the magnitude of the change in means ($\mu$) of the Gaussians, associating larger displacements with higher uncertainty. The motivation behind this lies in the fact that $\Sigma_{\hat{G}}$ can be approximated with the empirical Fisher Information matrix [43], $\Sigma_{\hat{G}}\approx\nabla_{G}\mathcal{L}(x_{k},y_{k})\nabla_{G}\mathcal{L}(x_{k},y_{k})^{\intercal}$, which in turn can be directly linked to the square of the updates of Gaussian parameters tuned by the gradient descent algorithm. In practice, we observe that the square of the magnitude of the updates outlined above is sensitive to measurement noise (in Sec. V-A), so we use the magnitude of the change instead.
@@
Empirically, we observe that the means of Gaussians located near boundaries between observed and unobserved space exhibit significant changes over successive updates, indicating high uncertainty, as visualized in Fig. 2. Consequently, directing exploration toward these high-uncertainty Gaussians is analogous to following frontiers in grid-based exploration, making this heuristic effective for autonomous exploration. Importantly, through careful bookkeeping, Gaussians that remain unchanged in the most recent update retain their previously computed displacement values, ensuring that our heuristic provides a consistent measure of each Gaussian’s movement over time. In addition to the boundaries between observed and unobserved space, areas within observed space that have insufficient measurements also exhibit relatively high uncertainty. This further directs the robot to explore these areas and gather more information, thereby improving the overall map quality.

### v3

@@ Abstract
We propose a framework for active mapping and exploration that leverages Gaussian splatting for constructing dense maps. Further, we develop a GPU-accelerated motion planning algorithm that can exploit the Gaussian map for real-time navigation. The Gaussian map constructed onboard the robot is optimized for both photometric and geometric quality while enabling real-time situational awareness for autonomy. We show through viewpoint selection experiments that our method yields comparable Peak Signal-to-Noise Ratio (PSNR) and similar reconstruction error to state-of-the-art approaches, while being orders of magnitude faster to compute. In closed-loop physics-based simulation and real-world experiments, our algorithm achieves better map quality (at least 0.8dB higher PSNR and more than 16% higher geometric reconstruction accuracy) than maps constructed by a state-of-the-art method, enabling semantic segmentation using off-the-shelf open-set models. Experiment videos and more details can be found on our project page: https://tyuezhan.github.io/RT_GuIDE/
@@ I Introduction
Active mapping is a problem of optimizing the trajectory of an autonomous robot in an unknown environment to construct an informative map in real-time. It is a critical component of numerous real-world applications such as precision agriculture, infrastructure inspection, and search and rescue missions. While nearly all tasks rely on recovering accurate metric information to enable path planning, many also require more fine-grained information. Recent advances in learned map representations from the computer vision and graphics communities [1, 2] have opened up new possibilities for active mapping and exploration while maintaining both geometrically and visually accurate digital twins of the environment.
While prior work effectively solves the problem of information-driven exploration [3, 4] or frontier-based exploration [5, 6], in this work we consider the additional problem of generating radiance fields while also performing autonomous navigation. Prior work has also proposed information metrics using novel learned scene representations that are capable of high quality visual reconstruction but are incapable of running in real-time onboard a robot [7]. To enable efficient mapping and planning in these novel representations, we consider approximation techniques for computing the information gain. Further, we consider the problem of generating high-quality maps that are capable of novel-view synthesis that can be used for downstream applications.
Fig. 1 shows the elements of our approach. We use the Gaussian splatting approach proposed in SplaTAM [8] to generate maps. We propose an efficient information gain metric and use a hierarchical planning framework to plan high-level navigation targets that yield maximal information in the environment and low-level paths that are dynamically feasible, collision-free and maximize the local information of the path. We utilize 3DGS as a unified representation for both mapping and planning. We avoid the cost of maintaining additional volumetric representations – achieving the same granularity of a 3DGS map with a voxel map would require around 19.2 GB of RAM for a $40m\times 40m\times 3m$ space at 1 $cm$ resolution. With our GPU-based trajectory planner, we achieve an 18x speedup compared to a CPU-based implementation, as detailed in Sec. V-B. This allows us to generate collision-free trajectories on millions of Gaussians efficiently. Our system runs real-time onboard a fully autonomous unmanned ground vehicle (UGV) to explore an unknown environment while generating a high-fidelity visual representation of the environment.
@@
    A unified framework for online mapping and planning built only on Gaussian splatting, eliminating the need for maintaining additional volumetric representations.
@@ II Related Work
Map Representation. To effectively construct a map of the environment, numerous map representations have been proposed in the robotics community. The most intuitive but effective volumetric representation has been widely used. Voxel-based representation could maintain information such as occupancy [9] or signed distance [10]. With the recent application of semantic segmentation, semantic maps that contain additional task-relevant information have been proposed [11, 12, 13]. With the recent advances in learned map representations in the computer vision community, Neural Radiance Fields (NeRF) [1] and 3D Gaussian Splatting (3DGS) [2] have become popular representations for robotic motion planning [7, 14, 15]. In this work, we study the problem of active mapping with the 3DGS representation.
@@
Active Mapping. The problem of exploration and active Simultaneous Localization and Mapping (SLAM) has been widely studied in the past decade. The classical exploration framework uses a model-based approach to actively navigate towards frontiers [5] or waypoints that have the highest information gain [3, 4, 7]. Some recent work combines the idea of frontier exploration and the information-driven approach to further improve efficiency [16, 17, 18]. However, most of the existing work developed their approaches based on classical map representations. In this work, we instead consider the active mapping problem with a 3DGS representation. Bayesian neural networks [19, 20] and deep ensembles [7] are common approaches for estimating uncertainty in learned representations. Radiance field representations provide additional possibilities for estimating uncertainty through the volumetric rendering process [21]. NARUTO [22] uses an uncertainty-learning module to quantify uncertainty for active reconstruction. CG-SLAM [23] uses the difference between rendered depth and observations, and the alignment of $\alpha$-blended and median depth as measures of uncertainty. FisherRF [24] leverages Fisher information to compute pixel-wise uncertainty on rendered images. Where prior work in implicit and radiance field representations use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a map represented by Gaussians encodes physical parameters of the scene, which motivates estimating information from the Gaussian parameters directly without rendering.
@@
Planning in Radiance Fields. Prior work has also considered planning directly in radiance fields. GS-planner [15] plans trajectories in a Gaussian map and uses observability coverage and reconstruction loss stored in a voxel grid as an approximation of information gain for exploration. Sim-to-real approaches leverage radiance fields as a simulator for imitation learning [25]. Adamkiewicz et al. [26] utilize the geometric fidelity of the representation to first map the environment and then perform trajectory optimization to plan paths in these environments. Splat-nav [14] uses a pre-generated Gaussian map to compute safe polytopes to generate paths. The probabilistic representations of free space has also been utilized for motion planning [27] where authors use uncertainties in the learned representation to provide guarantees on collision-free paths. While prior work considers planning in radiance fields, they typically maintain a separate volumetric representation for planning which requires additional memory usage onboard the robot. On the other hand, maintaining a secondary representation may lead to inconsistencies in collision checks for planning. Existing work that performs collision checking with 3DGS requires subsampling to provide soft constraints. In this work, our proposed planning module performs dense checks directly against all Gaussians for collision avoidance.
@@ III Problem Specification & Preliminaries
Figure 2: The proposed active mapping framework. The proposed framework contains two major components, the planning module and the mapping module. As can be seen in the figure, the Mapping module ([A]) takes in RGB, depth and pose measurements, and updates the map representation at every step and computes the utility of cuboidal regions (Sec. IV-B1). The utility of each region is then passed to the planning module which comprises the topological graph and motion primitive library ([B]). The trajectory planner in turn attempts to plan a path to goal that maximizes information gathering (queried from the mapper). The planned trajectory is executed by the robot to get a new set of observations.
@@
$\mathop{\arg\min}\limits_{\hat{G},\ x_{1:T}}$ $)\displaystyle\sum_{{test}\in\mathcal{T}}\mathcal{L}(y^{*}_{test},h(x^{*}_{test},\hat{G}$)) (1)
@@
$).\Phi(x_{1:k},y_{1:k}):=\mathop{\arg\min}\limits_{\hat{G}}\sum_{s=1}^{k}\mathcal{L}(y_{s},h(x_{s},\hat{G}$)). (2)
Given the posterior estimate of the map $\hat{G}$, we find the subsequent viewpoint $x_{k+1}$ that maximizes the information gain $\mathcal{I}(\cdot)$ given by the mutual information between the map $\hat{G}$ and the measurement $y_{k+1}$ at the corresponding pose:
$\max_{x_{k+1}}\ \mathcal{I}(\hat{G};y_{k+1}\mid x_{k+1}).$ (3)
@@
For a given pixel $p$ in the image, its color $\mathcal{C(\cdot)}$ can be obtained as
[…]
Similarly, the depth $\mathcal{D(\cdot)}$ can be obtained by replacing $c_{i}$ with $d_{i}$. At every iteration, we render color image $\hat{y}_{c}$ and depth image $\hat{y}_{d}$. We set the function $\mathcal{L}$ (from eq. (2)) to
$\mathcal{L}=\frac{1}{N_{p}}\sum_{p\in y}\left(|\hat{y}_{d}-y_{d}|+\lambda_{1}|\hat{y}_{c}-y_{c}|\right)+\lambda_{2}(1-SSIM(\hat{y}_{c},y_{c}))$
@@ IV Method
Our proposed framework comprises a mapping module and a planning module, as illustrated in Fig. 2. The mapping module (Sec. IV-A) accumulates measurements and poses to reconstruct the environment and computes the uncertainty of Gaussians in the map. This is then passed to the planning module for planning guidance paths (Sec. IV-B1) and trajectories (Sec. IV-B2).
@@ IV-A Mapping & Uncertainty Estimation
We build the 3DGS mapping module upon SplaTAM [8] with isotropic Gaussians. As in [8], we set $\lambda_{1}=0.4,\lambda_{2}=0.1$. At the beginning of each mapping iteration, we initialize new Gaussians at positions corresponding to the current color and depth measurements. The parameters of these Gaussians are refined through backpropagation. As in [2], we prune Gaussians with low opacity or large radius.
@@
Uncertainty Estimation. The objective for finding the next-best-view can then be expressed as maximizing the mutual information
[…]
where ${H}(\cdot)$ is the Shannon entropy. We have
$H(y_{k+1}\mid x_{k+1})=-\int p(y_{k+1}\mid x_{k+1})\log p(y_{k+1}\mid x_{k+1}))dy_{k+1}$
with $p(y_{k+1}\mid x_{k+1})=\int p(y_{k+1}\mid\hat{G},x_{k+1})p(\hat{G})d\hat{G}$ and
$H(y_{k+1}\mid$
[…]
$\biggl[\int p(y_{k+1}\mid\hat{G},x_{k+1})\log p(y_{k+1}\mid\hat{G},x_{k+1})\,dy_{k+1}\biggr]d\hat{G}$
@@
We make the following assumptions in this work to approximate the information gain efficiently: (a) $\hat{G}\sim\mathcal{N}(\mathrm{E}[G],\mathrm{Var}[G])$, where $G$ is the random variable that denotes the map; (b) $p(y_{k+1}\mid\hat{G})=\mathcal{N}(h(x_{k+1},\hat{G}),\Sigma_{y})$; with (c) constant isotropic variance $\Sigma_{y}$. For the marginal entropy ${H}(y_{k+1}\mid x_{k+1})$, given assumptions (a) and (b), and linearization of $h(\cdot)$ around $\mathrm{E}[G]$, we can approximate the distribution $p(y\mid x)=\mathcal{N}(h(x,\mathrm{E}[G]),J_{k+1}\mathrm{Var}[G]J_{k+1}^{\intercal})$, where $J_{k+1}=\partial h(x_{k+1},\hat{G})/\partial G$. Under assumptions (b) and (c), the conditional entropy does not depend on $G$. Recall the entropy of a $d_{Z}$ dimensional Gaussian distribution $\mathcal{N}(\mu_{Z},\Sigma_{Z})$ is $\frac{d_{Z}}{2}\log(2\pi e)+\frac{1}{2}\log\det(\Sigma_{Z})$. Dropping the constant terms, the information gain can be simplified to
[…]
@@
Using the first order approximation of $\log\det$ about the identity, we can define a proxy metric for scoring next best view as
[…]
@@
As in [24], assumption (c) allows us to simplify the optimization by ignoring the effect of $\Sigma^{-1}_{y}$. Evaluating $J_{k+1}$ is computationally expensive because it typically requires rendering at new poses. To achieve real-time evaluation of multiple candidate viewpoints, we replace the Jacobian term $J_{k+1}$ with a binary matrix that corresponds to its sparsity pattern. Intuitively, this captures the set of Gaussians within the field of view at pose $x_{k+1}$. The key insight behind these approximations is that we estimate information gain using the current uncertainty, based on the assumption that local measurements lead to precise local map estimates. We show empirically that the exploration performance based on our metric (in Sec. V-A2) is similar to the performance based on eq. (7). We propose a heuristic for estimating uncertainty $\mathrm{Var}[G]$ based on the magnitude of the change in means ($\mu$) of the Gaussians, associating larger displacements with higher uncertainty. The motivation behind this is that $\mathrm{Var}[G]$ can be approximated with the empirical Fisher Information matrix [28], $\mathrm{Var}[G]\approx\nabla_{G}\mathcal{L}(x_{k},y_{k})\nabla_{G}\mathcal{L}(x_{k},y_{k})^{\intercal}$, which in turn can be directly linked to the square of the updates of Gaussian parameters tuned by the gradient descent algorithm. In practice, we observe that the square of the magnitude of the updates outlined above is sensitive to measurement noise (in Sec. V-A3), so we use the magnitude of the change instead.
@@
Empirically, we observe that the means of Gaussians located near boundaries between observed and unobserved space exhibit significant changes over successive updates, indicating high uncertainty, as visualized in Fig. 2. Consequently, directing exploration toward these high-uncertainty Gaussians is analogous to following frontiers in grid-based exploration, making this heuristic effective for autonomous exploration. Through careful bookkeeping, Gaussians that remain unchanged in the most recent update retain their previously computed displacement values, ensuring that our heuristic provides a consistent measure of each Gaussian’s movement over time. In addition to the boundaries between observed and unobserved space, areas within observed space that have insufficient measurements also exhibit relatively high uncertainty. This further directs the robot to explore these areas and gather more information, improving the overall map quality.